OpenStudy (anonymous):

Write the system {2y+7z= -3, -3x-3y=3, 7x-8z=3} as a matrix equation @satellite73 @phi @thomaster Someone please help?????

OpenStudy (phi):

list your variables in alphabetical order: x,y,z and make them into a vector \[\left[\begin{matrix}x \\ y \\ z\end{matrix}\right]\]

OpenStudy (p0sitr0n):

x y z c 0 2 7 -3 -3 -3 0 3 7 0 -8 3

OpenStudy (phi):

now put a matrix on the left side of that vector. each row of the matrix will be one equation. It will multiply the vector of variables For example 2y+7z= -3 notice that this is the same as 0*x + 2*y +7*z = -3 the row of the matrix will 0 2 7

OpenStudy (anonymous):

okay thank you so much!! I think I get it now :)

OpenStudy (phi):

on the other side of the equal sign will be a vector of the "constants"

OpenStudy (phi):

\[\left[\begin{matrix}0 & 2 & 7 \\ ? & ? & ? \\ ? & ? & ?\end{matrix}\right]\left[\begin{matrix}x \\ y \\ z\end{matrix}\right]= \left[\begin{matrix}-3 \\ ? \\ ?\end{matrix}\right]\]

OpenStudy (anonymous):

[0 2 7] [x] [-3] [-3 -3 0] [y] = [3] [7 0 -8] [z] [3] ??? is this right?

OpenStudy (phi):

yes, looks good.

OpenStudy (phi):

If you know how to multiply a matrix times a vector, you can "re-create" the equations on the other hand, once you have a matrix form, you can use all the tools of linear algebra to solve these equations.

OpenStudy (anonymous):

Thank you so much! I appreciate all your help